at a restaurant, the bill for two coffees and 3 hot chocolates as 10.95. At another table the bill for one coffee and 2 hot chocolates was 6.65. How much did each type of beverage cost?

I really appreciate your help. I don't even know where to begin.

Answer by mmm4444bot(95)

Hello There:
Begin by assigning variable names for the two unknown amounts.
x = the price of a coffee
y = the price of a hot chocolate
From the given information, 2 coffees and 3 hot chocolates totaled $10.95. This alows us to write an equation:
2*x + 3*y = 10.95
Likewise, the other information leads to:
x + 2*y = 6.65
Solve the second equation for x:
x = 6.65 - 2*y
Substitute this expression for x in the first equation:
2*(6.65 - 2*y) + 3*y = 10.95
Can you now solve this equation for y? Once you know the value of y, you can find x because:
x = 6.65 - 2*y
~ Mark